Diameters of bolts produced by a particular machine are normally distributed




with mean 0.760 cm and standard deviation 0.012 cm. Specifications call for




diameters from 0.720 cm to 0.780 cm.





1
Expert's answer
2022-04-21T16:14:33-0400

"P(0.720<X<0.780)=P(\frac{0.720-0.760}{0.012}<Z<\frac{0.780-0.760}{0.012})=P(-3.33<Z<1.66)=0.9515-0.00043=0.95107"